OpenStudy (anonymous):

What is the value of c? –9c = 63 c =

OpenStudy (anonymous):

-7

sammixboo (sammixboo):

All you need to do is divide -9 on both sides of the equation \(\color{#20bd23}{-9c~=~63}\) \(\color{#20bd23}{\dfrac{-9c}{-9}~=~\dfrac{63}{-9}}\) \(\color{#20bd23}{c~=~\dfrac{63}{-9}}\) So what is \(\color{#20bd23}{63\div-9}\)?

OpenStudy (anonymous):

C=-7

OpenStudy (anonymous):

7

sammixboo (sammixboo):

Please do not give out the direct answer :) Thanks!

sammixboo (sammixboo):

Ah, try again, Asker!

OpenStudy (anonymous):

oh i didnt see the negative 9

sammixboo (sammixboo):

Ah, it's okay!

OpenStudy (anonymous):

-7

sammixboo (sammixboo):

Correcto!

OpenStudy (anonymous):

YAY!

sammixboo (sammixboo):

If you need anymore help, tag me \(\color{#20bd23}{@sammixboo}\)